Find the limit f(x),if it exists.sketch the graph of f(x).

Limit x approaching zero f(x),where f(x)=2-x,x less than or equal zero,. x+2, x>0


1
Expert's answer
2022-06-27T06:31:58-0400
"\lim\limits_{x\to0^-}f(x)=\lim\limits_{x\to0^-}(2-x)=2-0=2"

"\lim\limits_{x\to0^+}f(x)=\lim\limits_{x\to0^+}(x+2)=0+2=2"

Since "\lim\limits_{x\to0^-}f(x)=2=\lim\limits_{x\to0^+}f(x)," then "\lim\limits_{x\to0}f(x)" exists, and


"\lim\limits_{x\to0}f(x)=2"